OpenStudy (anonymous):

Nancy bought 360 tropical fish for a museum display. She bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of angelfish Nancy bought and y represents the number of goldfish she bought?

OpenStudy (anonymous):

A. x – y = 360 y = 8x

OpenStudy (anonymous):

B. x + y = 360 y = 8x

OpenStudy (anonymous):

C. x + y = 8 y = 360x

OpenStudy (anonymous):

D. x + 8y = 360 y = 8x

OpenStudy (anonymous):

@CoconutJJ

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (solomonzelman):

\(\large\color{blue}{ y }\) is: number of \(\normalsize\color{black}{\rm goldfish }\) \(\large\color{blue}{x }\) is: number of \(\normalsize\color{black}{\rm angelfish }\)

OpenStudy (solomonzelman):

you have: `360 tropical fish for a museum display.` (saying that \(\normalsize\color{black}{\rm goldfish }\) and \(\normalsize\color{black}{\rm angelfish }\) altogether is 360 fishes) `She bought 8 times as many goldfish as angelfish. ` (saying that the number of \(\normalsize\color{black}{\rm goldfish }\) is \(\normalsize\color{blue}{\rm8~times~greater }\) than the number of \(\normalsize\color{black}{\rm angelfish }\))

OpenStudy (solomonzelman):

Can you write your equations with all I gave you?

OpenStudy (anonymous):

so is it B ?

OpenStudy (anonymous):

@demonchild99